OpenStudy (anonymous):

A road sign is in the shape of a regular heptagon. What is the measure of each angle on the sign? Round to the nearest tenth. 128.6 900 64.3 231.4

OpenStudy (anonymous):

A

OpenStudy (anonymous):

I think the multitude of questions from kamuela means there is very little understanding is taking place. Perhaps K needs to ask specific questions on how to solve the problem rather than getting answers for homework.

OpenStudy (anonymous):

MrYantho - you have a good point. However, it is up to the people to do effort part.

OpenStudy (anonymous):

Sum of all angles of a polygon is governed by the relation S = 180*(n-2) where n is the no of sides in the polygon (n greater than or equal to 3) for heptagon n = 7. therefore each angle measure = 180*(n-2)/n = 180*(7-2)/7 = 180*5/7 = 900/7 = 128.57142857... therefore approximately 128.6. Thus option A.